OpenStudy (mackenzie2013):

Hayden invested $750 for three years at an interest rate of 0.035. How much interest did he earn in that time?

OpenStudy (mackenzie2013):

78.75?

OpenStudy (anonymous):

Yep :)

OpenStudy (kamibug):

Sorry, but @NerdyMackenzie missed an initial step. Actually, first you need to see how much .035 is of 750 (because interst rates are percentages) and then add it to 750 and then multiply that by 3 years. So the proper steps would be... .035 * 750 = 26.25 750 + 26.25 = 776.25 776.25 * 3 = 2328.75 $2328.75 is the amount of money he would have after three years including the interests. =D Hope this helps!

OpenStudy (anonymous):

I am pretty sure my answer is right so..

OpenStudy (kamibug):

Google it. xD

OpenStudy (anonymous):

(750.00)(0.035.)(3) $73.125